The company’s year end is September 30. Prepare the journal entries required at each date, as well as any necessary adjusting entries required on Sept. 30 at the year end.  

April 1

Investors contributed $350,000 in cash to start the Company.

April 15

The Company purchased a warehouse for $125,000. Depreciation for the six months totaled $1,000.

April 17

The Company purchased sixteen golf bags for $8,000 total ($500 each). The company paid for half of the stock with cash and was billed for the remainder.

May 1

The company paid $6,000 for advertisements that will run monthly May 2019 through May 2020.

July 1

The Company borrowed $200,000 at 4% interest rate from Giant Bank; the principal will be due in five years. The first interest payment will be due July 1, 2020.

Aug. 20

The company sent a $3,500 payment to its supplier for the golf bags it purchased in April.

Aug. 21

A customer prepaid $10,000 for a special order that will be delivered by in January.

Sept. 15

The company declared and paid a dividend of $500.

Sept. 30

The company sold 12 golf bags during the last six months for $1,500 each; however, the company only collected $3,000 in cash; the remaining amount was billed to creditworthy customers and is expected to be paid in February 2020.

Sept. 30

Income taxes totaled $2,000. Assume that taxes will be paid in February 2020.
Answer #1
Journal Entries for the transaction given:
Date Particulars Debit Credit
1st April Cash $350,000
Owners Capital $350,000
(To record contribution made for cash )
15th April Warehouse $125,000
Cash $125,000
(To record warehouse purchased for cash )
17th April Inventory $8,000
Cash $4,000
Accounts payable $4,000
(To record purchases made)
(Assuming Perpectual method is followed)
1st May Prepaid advertisement $6,000
Cash $6,000
(To record advertisement paid for cash )
1st July Cash $200,000
Gaint Bank Loan $200,000
(To record amount borrowed from bank)
20th Aug Accounts Payable $3,500
Cash $3,500
(To record cash paid to supplier)
21st Aug Cash $10,000
Unearned Revenue $10,000
(Being cash received for sales to be made
in January)
15th Sept Dividend $500
Cash $500
(To record dividend paid)
30th Sept Cash $3,000
Accounts Receivable $15,000
Sales Revenue $18,000
Sales revenue = 12*1500 =18000
(To record Sales made on cash and credit
30th Sept Income Tax Expense $2,000
Income Tax Expense Payable $2,000
(To record taxes payable in Febuary 2020)
Adjusting Entries:
30th Sept Depreciation $1,000
Accumalated Depreciation $1,000
(To record depreciation)
30th Sept Advertisement $2,500
Prepaid Advertisement $2,500
(To record advertisement for the period)
$6000 *5Months /12 Months = $2500
5 months i.e from may - sept
30th Sept Interest Expense $2,000
Interest Expense Payable $2,000
(To record Interest expense till september end )
(Loan amount * Rate of Interest * Months in a year /12
$200000*4%*3(july- sep)/12